Assume you have a 36.0 Ω, a 50.0 Ω, and a 700 Ω resistor.

a. What is the smallest effective resistance you could obtain by connecting these three resistors together?

b. How would you have to connect the resistors to obtain the smallest effective resistance?

c. What is the largest effective resistance you could obtain by connecting these three resistors together?

d. How would you have to connect the resistors to obtain the largest effective resistance?
Answer #1

a) Req= 36x50x700/(36x50+50x700+700x36) = 20.322 ohm
b) all in parallel
c) Rmax= 36+50+700= 786 ohm
d) all in series